# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v1

@@ I Introduction
Reliable quantum communication is a critical area of research, essential for scaling quantum systems, enabling the exchange of quantum information across distances, and facilitating multiparty protocols. Quantum states are inherently fragile, requiring the implementation of quantum error correction to mitigate the effects of noise and decoherence [2, 3]. In this work, we explore advanced decoding techniques for quantum error-correcting codes to enhance the reliability and efficiency of quantum communication.
@@
The decoding problem for quantum codes with high degeneracy is particularly challenging for BP, as seen in topological codes [22, 23, 24, 25, 26, 27, 28]. To address these challenges, several remedies for the BP algorithm have been introduced, including techniques such as random perturbation [8], enhanced feedback [9, 10], check-matrix augmentation [11], message normalization and offset [12, 13], and trapping set analysis [29, 30]. These methods aim to improve the decoding accuracy and convergence of the BP algorithm. A recently developed modified BP4 algorithm, known as MBP4, incorporates additional memory effects and has demonstrated the ability to handle topological codes [31]. The adaptive version of MBP4, termed AMBP4, further improves BP performance while maintaining nearly linear time complexity [31]. Nonetheless, its practical application involves an unknown parameter search and may not be fully parallelized.
We investigate ordered statistic decoding (OSD) in quantum coding theory as a complementary approach to BP. Fossorier and Lin [32] first proposed OSD in classical coding theory as an approach to approximate maximum-likelihood decoding [33, 34, 35, 36, 37, 38, 39, 40, 41]. The OSD algorithm searches for possible solutions by sorting the reliability measures of symbols while maintaining controllable complexity [32, 42, 43, 44, 45]. In cases where BP fails to find a valid error for a given syndrome, syndrome-based OSD can be utilized [46]. Moreover, OSD has been demonstrated to be particularly useful for degenerate quantum codes by Panteleev and Kalachev [42].
The core function of OSD is Gaussian elimination on the parity-check matrix of dimension $m\times n$, which identifies the independent unreliable bits and has a complexity of ${\cal O}(nm^{2})$. The remaining reliable bits can be used to regenerate the most probable error. In an order-$w$ OSD, the set of reliable bits can be flipped up to $w$ bits, generating a list of probable solutions. However, this technique incurs at least a cubic time complexity overhead from Gaussian elimination, along with additional costs for high-order examination, despite its effectiveness.
Fossorier et al[46] identified specific candidate pruning conditions to reduce ineffective high-order OSD processing, as such operations do not yield improved error candidates. Many techniques have been proposed in classical coding theory to reduce the complexity of high-order OSD, by focusing on more probable candidates, thereby reducing the number of OSD candidates required [47, 48, 49, 50, 51, 52].
@@
However, there is a great deal of room for improvement over the current postprocessing techniques in the literature under the code capacity noise model. These methods primarily rely on binary BP to handle quaternary errors, neglecting error correlations, which suggests that binary decoders are inherently suboptimal. In contrast, the quaternary AMBP4 decoder in [31] has been shown to outperform all the aforementioned decoders without requiring postprocessing. For supporting examples, please refer to Figure 2, Figure 4, and Figure 7 in the simulation section.
Since OSD acts as a complementary strategy to BP for cases where BP fails, BP-OSD is naturally expected to surpass the performance of AMBP4. In this paper, we propose a BP4-based OSD to achieve this improvement. Additionally, we introduce an efficient OSD algorithm tailored for quantum codes.
@@
Next, we consider approximate degenerate OSD in quantum codes. A typical OSD generates a list of error candidates for a given error syndrome and selects the one with the minimum weight, which is optimal for classical codes. For quantum codes, however, the optimal decoding criterion involves choosing the error coset with the highest probability, a criterion that is challenging to implement in practice. To address this, we consider approximate degenerate decoding in OSD by focusing only on relatively low-weight terms within an error coset.
@@
Additionally, we introduce a degenerate error candidate pruning condition for stabilizer codes. We show that certain bit-flips in OSD correspond to multiplying stabilizers, resulting to degenerate error candidates. These bit-flips may alter the error weight but do not change the error coset, allowing them to be ignored in approximate degenerate OSD. For improving OSD efficiency, using a minimum weight decoding criterion is beneficial. A key observation is that if all bit-flips correspond to multiplying a stabilizer, there is no need to implement high-order OSD. We will demonstrate that this is indeed the case at low error rates by simulations in Section VI-G. We note that a degeneracy condition is also utilized in ambiguity clustering to facilitate the resolution of subproblems [56].
@@
Integrating all the mentioned techniques, we propose an approximate degenerate OSD algorithm, abbreviated as ADOSD4. We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [55], rotated toric codes and rotated surface codes [24], (6.6.6) and (4.8.8) color codes [27], twisted XZZX codes on a torus [58], generalized hypergraph-product (GHP) codes [42, 43], and lift-connected surface (LCS) codes [59].
@@
The results demonstrate that our proposed schemes outperform previous approaches in the literature, both in terms of error thresholds for topological codes and low error rate performance (logical error rate around $10^{-6}$). Both our MBP4+OSD4 and MBP4+ADOSD4 decoders outperform other binary BP-OSD schemes in these quantum codes. We achieve thresholds of approximately 17.5%–17.7% for the toric, surface, and twisted XZZX codes, and thresholds of 15%–15.42% for the (4.8.8) and (6.6.6) color codes, as summarized in Table II. Comparisons of our results with existing threshold values for the toric codes are provided in Table III. We also compare the performance of various reliability metrics on the GHP code, justifying our choice of reliability metric.
@@
Finally, we analyze the time efficiency of our OSD schemes for various topological codes of comparable distances to the $[[882,48,16]]$ GHP code using single-threaded CPU implementations. The simulation results, summarized in Table IV, show that the time consumption of ADOSD4 is approximately 1% to 3% of that for order-2 OSD4 with similar decoding performance. Moreover, Table VI presents the time consumption of ADOSD4 in terms of BP iterations for surface codes with distances from 11 to 15. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 2 to 3 BP4 iterations, whereas a successful BP trial typically takes 2 to 3 iterations on average. This suggests that ADOSD4 is an excellent complement to BP, especially since ADOSD4 is used only when BP fails.
@@
This paper is organized as follows. We introduce quantum stabilizer codes in Section II. Section III discusses reliability metrics and the associated OSD algorithm. The reliable subset reduction method is presented in Section IV. We then cover approximate degenerate decoding in Section V, and provide simulation results in Section VI. Finally, we conclude in Sec. VII.
@@ II Quantum stabilizer codes
Let $I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$ $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$ $Z=\begin{bmatrix}1&0\\ 0&-11\end{bmatrix},$ and $Y=iXZ$ denote the Pauli matrices. Consider the $n$-fold Pauli group
$\mathcal{G}_{n}\triangleq\left\{cM_{1}\otimes\dots\otimes M_{n}:c\in\{\pm 1,\pm i\},M_{j}\in\{I,X,Y,Z\}\right\}.$
Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues ±$\pm 1$. Any two Pauli operators either commute or anticommute with each other. The weight of a Pauli operator $E\in{\cal G}_{n}$, denoted $\mathrm{wt}\left(E\right)$, refers to the number of its nonidentity components.
We consider Pauli errors in this paper, assuming independent depolarizing errors with rate $\epsilon$. Each qubit independently experiences an $X$, $Y$, or $Z$ error with probability $\epsilon/3$ and no error with probability $1-\epsilon$. Therefore, at a small error rate $\epsilon$, low-weight Pauli errors are more likely to occur.
@@
$\mathcal{C}(\mathcal{S})=\left\{\ket{\psi}\in\mathbb{C}^{2^{n}}:S\ket{\psi}=\ket{\psi},~{}\forall\,S\in\mathcal{S}\right\}.$
@@
An error $E\in{\cal G}_{n}$ can be detected by ${\cal C}({\cal S})$ through stabilizer measurements if $E$ anticommutes with some of its stabilizers. Suppose that $\{S_{i}\}_{i=1}^{m}$, where $m\geq n-k$, is a set of stabilizers that generates ${\cal S}$. Their binary measurement outcomes are referred to as the error syndrome of $E$. Let $N({\cal S})\subset{\cal G}_{n}$ denote the normalizer group of ${\cal S}$, which consists of the Pauli operators that commute with all stabilizers. Consequently, if an error is in $N({\cal S})$, it will have zero syndrome and cannot be detected. Note that if an error is a stabilizer, it has no effect on the code space. An element in the normalizer group $N({\cal S})$ that is not a stabilizer, up to a phase, is called a nontrivial logical operator as it changes the logical state of a code . Therefore, the minimum distance of ${\cal C}({\cal S})$ is defined as the minimum weight of a nontrivial logical operator. It can be observed that $E$ and $ES$ for $S\in{\cal S}$ have the same effects on the codespace. They are called degenerate to each other. We say that a quantum code is highly degenerate if it has many stabilizers with low weight relative to its minimum distance.
@@
$\sum_{S\in{\cal S},\mathrm{wt}\left(S\right)\leq\delta}\Pr\big{\{}\hat{E}S\big{\}}.$ (1)
@@ II-A Binary representations
Without loss of generality, we assume that each stabilizer in $\{S_{i}\}_{i=1}^{m}$ is of the form $S_{i}=S_{i1}\otimes S_{i2}\otimes\dots\otimes S_{in}$, where $S_{ij}\in\{I,X,Y,Z\}$. Then we can study the decoding problem in the binary vector space [4, 6], using a mapping $\varphi:$
@@
A check matrix of the stabilizer code ${\cal C}({\cal S})$ is a binary matrix ${\mathbf{S}}=[{\mathbf{S}}_{ij}]\in\{0,1\}^{m\times 2n},$ where
$\left[\begin{array}[]{c|c}{\mathbf{S}}_{ij}&{\mathbf{S}}_{i(j+n)}\end{array}\right]=\varphi(S_{ij}).$
An $n$-fold Pauli operator $E\in{\cal G}_{n}$ can also be represented by $\varphi(E)\triangleq{\mathbf{E}}=\left[\begin{array}[]{c|c}{\mathbf{E}}^{X}&{\mathbf{E}}^{Z}\end{array}\right]\in\{0,1\}^{2n}$, where ${\mathbf{E}}^{X}$ and ${\mathbf{E}}^{Z}$ are the indicator vectors of $X$ and $Z$ components of $E$, respectively. Then the error syndrome of $E$ is
${\mathbf{z}}={\mathbf{S}}\Lambda{\mathbf{E}}^{\top}\in\{0,1\}^{m\times 1},$ (2)
@@
Consequently, the decoding problem is to solve this system of linear equations with $2n$ binary variables ${\mathbf{E}}_{j}$ that are most probable. Note that ${\mathbf{S}}$ is of rank $n-k$ so the degree of freedom of this binary system is $n+k$.
@@ II-B Belief Propagation (BP) Decoding
Given a check matrix ${\mathbf{S}}$, a syndrome ${\mathbf{z}}$, and a depolarizing error rate $\epsilon$, BP performs the following steps:
@@
Steps 2) and 3) will be iterated for a maximum number of $T$ iterations, where $T$ is chosen in advance. If no valid estimate is obtained after $T$ iterations, BP will claim a failure.
Typically, the initial error distribution ${\mathbf{p}}_{i}$ is defined as $({1-\epsilon},\frac{\epsilon}{3},\frac{\epsilon}{3},\frac{\epsilon}{3})$. However, one may instead choose a fixed initialization by a given value $\epsilon_{0}\neq\epsilon$ for BP as long as BP performs well at this $\epsilon_{0}$. See more discussion on fixed initialization in [63, 13, 31].
@@ III Order-statistic decoding based on quaternary belief propagation
In the aforementioned syndrome decoding problem, we aim to solve a system of linear equations given by (2), for a given parity-check matrix ${\mathbf{S}}$ and an error syndrome ${\mathbf{z}}$ with $2n$ binary variables. The linear system has $n-k$ independent constraints, impliying that there are $n+k$ degrees of freedom in the vector ${\mathbf{E}}$. Given the values of these $n+k$ bits, one can reconstruct ${\mathbf{E}}$. If BP fails to provide a valid error estimate after $T$ iterations, OSD will be utilized.

### v2

@@ I Introduction
Reliable quantum communication is a critical area of research, essential for scaling quantum systems, enabling the exchange of quantum information across distances, and facilitating multiparty protocols. Quantum states are inherently fragile, requiring the implementation of quantum error correction to mitigate the effects of noise and decoherence [2, 3, 4].
@@
The decoding problem for quantum codes with high degeneracy is particularly challenging for BP, as seen in topological codes [23, 24, 25, 26, 27, 28, 29]. To address these challenges, several remedies for the BP algorithm have been introduced, including techniques such as random perturbation [9], enhanced feedback [10, 11], check-matrix augmentation [12], message normalization and offset [13, 14], and trapping set analysis [30, 31]. These methods aim to improve the decoding accuracy and convergence of the BP algorithm. A recently developed modified BP4 algorithm, known as MBP4, incorporates additional memory effects and has demonstrated the ability to handle topological codes [32].
We investigate ordered statistic decoding (OSD) as a complementary approach to BP in quantum coding theory. Originally proposed by Fossorier and Lin [33] in classical coding theory for approximating maximum-likelihood decoding [34, 35, 36, 37, 38, 39, 40, 41, 42], OSD sorts reliability measures of symbols to find probable solutions while maintaining controllable complexity [33, 43, 44, 45, 46]. Syndrome-based OSD has been used when BP fails to identify a valid error [47] and is particularly effective for degenerate quantum codes [43].
The core operation of OSD is Gaussian elimination on an $m\times n$ parity-check matrix, identifying unreliable bits with a complexity of $O(nm^{2})$. Order-$w$ OSD flips up to $w$ reliable bits to generate error candidates, but this incurs high computational costs. To mitigate complexity, Fossorier et al.[47] proposed candidate pruning, and various classical techniques further optimized high-order OSD by reducing the number of candidates examined[48, 49, 50, 51, 52, 53].
@@
However, there is a great deal of room for improvement over the current postprocessing techniques in the literature under the code capacity noise model. These methods primarily rely on binary BP to handle quaternary errors, neglecting error correlations, which suggests that binary decoders are inherently suboptimal. In contrast, the quaternary AMBP4 decoder (the adaptive version of MBP4) from [32] has been shown to outperform all the aforementioned decoders across various quantum codes, including topological codes [25], BB codes [56], and generalized hypergraph-product (GHP) codes [43, 44], without requiring postprocessing. For supporting examples, please refer to Figure 7 in Section VI.
OSD serves as a complementary strategy to BP, addressing cases where BP fails. As a result, BP-OSD is naturally expected to outperform AMBP4. In this paper, we propose a BP4-based OSD to achieve this improvement and introduce an efficient OSD algorithm for quantum codes.
@@
Next, we consider approximate degenerate OSD in quantum codes. A typical OSD generates a list of error candidates for a given error syndrome and selects the one with minimum weight, which is optimal for classical codes. For quantum codes, however, the optimal decoding criterion involves choosing the error coset with the highest probability, a criterion that is challenging to implement in practice. To address this, we consider approximate degenerate decoding in OSD by focusing only on relatively low-weight terms within a coset.
@@
Additionally, we introduce a degenerate error candidate pruning condition for stabilizer codes. We show that certain bit-flips in OSD correspond to multiplying stabilizers, resulting in degenerate error candidates. These bit-flips may alter the error weight but do not change the error coset, allowing them to be ignored in approximate degenerate OSD. To improve OSD efficiency, the minimum weight decoding criterion is applied. A key observation is that if all bit-flips correspond to multiplying a stabilizer, high-order OSD is unnecessary. We note that a degeneracy condition is also utilized in ambiguity clustering to facilitate subproblem resolution [57].
@@
Integrating all the mentioned techniques, we propose an approximate degenerate OSD algorithm, abbreviated as ADOSD4. We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [56], rotated toric codes and rotated surface codes [25], (6.6.6) and (4.8.8) color codes [28], twisted XZZX codes on a torus [59], GHP codes [43, 44], and lift-connected surface (LCS) codes [60].
@@
The results demonstrate that our proposed schemes outperform previous approaches in the literature, both in terms of error thresholds for topological codes and low error rate performance (logical error rate around $10^{-6}$). Our MBP4+OSD4 and MBP4+ADOSD4 decoders outperform other binary BP-OSD schemes in these quantum codes. We achieve thresholds of approximately 17.5%–17.7% for the toric, surface, and twisted XZZX codes, and thresholds of 15%–15.42% for the (4.8.8) and (6.6.6) color codes, as summarized in Table III. They even surpass the minimum weight perfect matching (MWPM) decoder on surface and toric codes [69, 70].
@@
Finally, we analyze the time efficiency of our OSD schemes for these quantum codes using single-threaded CPU implementations. The simulation results, summarized in Table IV, show that the time consumption of ADOSD4 for surface codes with distances from 11 to 15 is approximately 2% to 4% of that for order-2 OSD4 with similar decoding performance at logical error rate around $10^{-6}$. Moreover, the average time consumption per codeword for ADOSD4 is approximately equivalent to 2 to 3 BP4 iterations, whereas a successful BP trial typically takes 2 to 3 iterations on average. This suggests that ADOSD4 is an excellent complement to BP, especially since ADOSD4 is used only when BP fails.
@@
This paper is organized as follows. We introduce stabilizer codes in Section II. Section III discusses reliability metrics and the associated OSD algorithm. The reliable subset reduction method is presented in Section IV. We then cover approximate degenerate decoding in Section V, and provide simulation results in Section VI. Finally, we conclude in Sec. VII.
@@ II Quantum stabilizer codes
Let $I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$ $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$ $Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},$ and $Y=iXZ$ denote the Pauli matrices. The $n$-fold Pauli group, $\mathcal{G}_{n}$, is the group of $n$-qubit Pauli operators, defined as
$\mathcal{G}_{n}\triangleq\left\{\omega M_{1}\otimes\dots\otimes M_{n}:\omega\in\{\pm 1,\pm i\},M_{j}\in\{I,X,Y,Z\}\right\}.$
Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues ±$\pm 1$. Any two Pauli operators either commute or anticommute with each other. The weight of a Pauli operator $g\in{\cal G}_{n}$, denoted $\mathrm{wt}\left(g\right)$, refers to the number of its nonidentity components.
We consider Pauli errors, assuming independent depolarizing errors with rate $\epsilon$. Each qubit independently experiences an $X$, $Y$, or $Z$ error with probability $\epsilon/3$ and no error with probability $1-\epsilon$. Therefore, at a small error rate $\epsilon$, low-weight Pauli errors are more likely to occur across $n$ qubits.
@@
$\mathcal{C}(\mathcal{S})=\left\{\ket{\psi}\in\mathbb{C}^{2^{n}}:g\ket{\psi}=\ket{\psi},~{}\forall\,g\in\mathcal{S}\right\}.$
@@
An error $e\in{\cal G}_{n}$ can be detected by ${\cal C}({\cal S})$ through stabilizer measurements if $e$ anticommutes with some of the stabilizers. Suppose that $\{g_{i}\}_{i=1}^{m}$, where $m\geq n-k$, is a set of stabilizers that generates ${\cal S}$. Their binary measurement outcomes are referred to as the error syndrome of $e$. Let $N({\cal S})\subset{\cal G}_{n}$ denote the normalizer group of ${\cal S}$, which consists of the Pauli operators that commute with all stabilizers. Consequently, if an error is in $N({\cal S})$, it will have zero syndrome and cannot be detected. Note that if an error is a stabilizer, it has no effect on the code space. An element in the normalizer group $N({\cal S})$ that is not a stabilizer, up to a phase, is called a nontrivial logical operator as it changes the logical state of a code . Therefore, the minimum distance of ${\cal C}({\cal S})$ is defined as the minimum weight of a nontrivial logical operator. It can be observed that $e\in{\cal G}_{n}$ and $eg$ for $g\in{\cal S}$ have the same effects on the codespace. They are called degenerate to each other. We say that a quantum code is highly degenerate if it has many stabilizers with low weight relative to its minimum distance.
@@
$\sum_{g\in{\cal S},\mathrm{wt}\left(g\right)\leq\delta}\Pr\big{\{}\hat{e}g\big{\}}.$ (1)
@@ II-A Binary representations
We can study the decoding problem in the binary vector space [5, 7], using a mapping $\varphi:{\cal G}_{1}\rightarrow\{0,1\}^{2}$
@@
We extend the $\varphi$ to ${\cal G}_{n}$ as follows: For $g=M_{1}\otimes M_{2}\otimes\dots\otimes M_{n}\in{\cal G}_{n}$,
$\varphi(g)\triangleq{\mathbf{g}}=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}&{\mathbf{g}}^{Z}\end{array}\right]=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}_{1}\cdots{\mathbf{g}}^{X}_{n}&{\mathbf{g}}^{Z}_{1}\cdots{\mathbf{g}}^{Z}_{n}\end{array}\right],$
where ${\mathbf{g}}^{X}\in\{0,1\}^{n}$ and ${\mathbf{g}}^{Z}\in\{0,1\}^{n}$ are the indicator vectors of $X$ and $Z$ components of $g$, respectively, such that $\left[\begin{array}[]{c|c}{\mathbf{g}}_{j}^{X}&{\mathbf{g}}_{j}^{Z}\end{array}\right]=\varphi(M_{j}).$
A check matrix of ${\cal C}({\cal S})$ corresponding a set of stabilizer generators $\{g_{i}\}_{i=1}^{m}$ is an $m\times 2n$ binary matrix ${\mathbf{H}}$ with rows $\varphi(g_{1}),\dots,\varphi(g_{m})$.
An error $e\in{\cal G}_{n}$ can be represented by ${\mathbf{e}}=\varphi(e)$ and its error syndrome is
${\mathbf{s}}={\mathbf{H}}\Lambda{\mathbf{e}}^{\top}\in\{0,1\}^{m\times 1},$ (2)
@@
Suppose this code has $2k$ independent logical operators, $\bar{X}_{i}$ and $\bar{Z}_{j}$ for $i,j=1,\dots,k$, which, together with the stabilizers $\{g_{i}\}$, generate the normalizer group $N({\cal S})$. The $2k\times 2n$ binary logical matrix ${\mathbf{L}}$ is defined with rows $\varphi(\bar{X}_{1}),\dots,\varphi(\bar{X}_{k}),\varphi(\bar{Z}_{1}),\dots,\varphi(\bar{Z}_{k})$. Since stabilizers commute with logical operators, we have ${\mathbf{H}}\Lambda{\mathbf{L}}^{\top}=\mathbf{0}$.
The decoding problem is to solve the system of linear equations (2) with $2n$ binary variables in ${\mathbf{e}}\in\{0,1\}^{2n}$ that are most probable. Note that ${\mathbf{H}}$ is of rank $n-k$ so the degree of freedom of this binary system is $n+k$.
An error $e\in{\cal G}_{n}$ is classified as a logical error for a decoder if the decoder fails or outputs $\hat{e}\in{\cal G}_{n}$ such that
${\mathbf{H}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}=\mathbf{0}$ (3)
but
${\mathbf{L}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}\neq\mathbf{0}.$ (4)
@@ II-B Belief Propagation (BP) Decoding
Given a check matrix ${\mathbf{H}}$, a syndrome ${\mathbf{s}}$, and a depolarizing error rate $\epsilon$, BP performs the following steps:
@@
Steps 2) and 3) will be iterated for a maximum number of ${\mathrm{T}}$ iterations, where ${\mathrm{T}}$ is chosen in advance. If no valid estimate is obtained after ${\mathrm{T}}$ iterations, BP will claim a failure.
@@ III Order-statistic decoding based on quaternary belief propagation
In the aforementioned syndrome decoding problem, we aim to solve a system of linear equations given by (2), for a given parity-check matrix ${\mathbf{H}}$ and an error syndrome ${\mathbf{s}}$ with $2n$ binary variables in ${\mathbf{e}}$. The linear system has $n-k$ independent constraints, impliying that there are $n+k$ degrees of freedom in the vector ${\mathbf{e}}$. Given the values of these $n+k$ bits, one can reconstruct ${\mathbf{e}}$. If BP fails to provide a valid error estimate after ${\mathrm{T}}$ iterations, OSD will be utilized.